OpenStudy (anonymous):

Grade 8 math will medal for best answer what is: -26 + 2x = -2

OpenStudy (anonymous):

-26 + 2x = -2 x=12

OpenStudy (anonymous):

2x = 24 x = 12

OpenStudy (anonymous):

sweetness thanks!

OpenStudy (dahlindubs11):

-24=-2x.... then divide each side by -2... you get +12

OpenStudy (anonymous):

Why do people always turn X into a negative? @dahlindubs11

OpenStudy (dahlindubs11):

ur not turning x into a negative your turning (2x) into a negative.... explanation is this... start with -26+2x=-2... add 2 to each side and subtract -2x from each side so you have numbers on one side and variables on other... -26+2=-24 and -2+2=0... that would make the equation -24+2x=0... then subtract the 2x from each side. 2x-2x=0 and 0-2x=-2x... now ur equation is -24=-2x. make sense?

OpenStudy (anonymous):

"...and subtract -2x from each side..." That Is what I mean.

OpenStudy (anonymous):

why not just 2x = 26 -2

OpenStudy (anonymous):

@dahlindubs11

OpenStudy (dahlindubs11):

that is also correct @Holt but I was trying to explain the basics of it...

OpenStudy (anonymous):

Bleh, when I was learning it, that was the biggest reason I always did so bad. Extra steps = more chances to mess up.